OpenStudy (anonymous):

if f(x)=3x^2 and g(x)=4x^3+1 what is the degree of (f*g)(x)

OpenStudy (solomonzelman):

is it fºg or f•g ?

OpenStudy (anonymous):

f * g

OpenStudy (solomonzelman):

that is f•g, right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Large\color{red}{ \displaystyle f(x)=3x^2 }\) \(\Large\color{blue}{ \displaystyle g(x)=4x^3+1 }\) -------------------------------- \(\Large\color{black}{ \displaystyle \left(\color{red}{f}~\cdot~\color{blue}{g}\right)(x)=\left(\color{red}{3x^2}\right)~\cdot~\left(\color{blue}{4x^3+1}\right) }\)

OpenStudy (solomonzelman):

the first term will give you the degree, you don't to know the rest of the actual function

OpenStudy (solomonzelman):

what is (3x\(^2\) • 4x\(^3\)) ? that will give you x raised to what power?

OpenStudy (anonymous):

12x^5

OpenStudy (solomonzelman):

yes, that is x raised to the 5th power (and that will be the biggest power you will have in your new function -- is that correct?) so what is the degree of your new function?

OpenStudy (solomonzelman):

Note: Degree - [means] the biggest power.

OpenStudy (anonymous):

5?

OpenStudy (solomonzelman):

Yes, there you go!

OpenStudy (solomonzelman):

That is it, you are done

OpenStudy (anonymous):

thank you so much the colors were very helpful!!!!

OpenStudy (solomonzelman):

You are welcome:)